# On the numerical stability of sketched GMRES

### v1

@@ 7 Conclusions
In this paper, we performed a backward stability analysis of sketched GMRES and showed how the backward error is largely influenced by the condition number of the generated Krylov basis. However, in [7, Figure 9], it was demonstrated that the backward error of sGMRES can still decrease even when the basis becomes ill-conditioned. To provide more insight into this finding from [7], we proposed a sharper bound that relies on $\|\hat{B}_{1:k}\rVert\|\hat{y}^{(k)}\rVert/\|\hat{x}^{(k)}\rVert$ rather than $\kappa(\hat{B}_{1:k})$. We then showed how restarting can be used to stabilize a sketched GMRES implementation, and proposed an adaptive restarting strategy for sGMRES based on our analysis. Our numerical experiments indicate that the convergence of sketched GMRES with our adaptive restarting approach often performs comparably to standard GMRES and has the potential to result in substantial time savings.
@@ Acknowledgments
All authors are supported by the Charles University Research Centre program No. UNCE/24/SCI/005 and by the European Union (ERC, inEXASCALE, 101075632). Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.

### v2

@@ 7 Conclusions
In this paper, we performed a backward stability analysis of sketched GMRES and showed how the backward error is largely influenced by the condition number of the generated Krylov basis. However, in [guttel2024sketch, Figure 9], it was demonstrated that the backward error of sGMRES can still decrease even when the basis becomes ill-conditioned. To provide more insight into this finding from [guttel2024sketch], we proposed a sharper bound that relies on $\|\hat{B}_{1:k}\rVert\|\hat{y}^{(k)}\rVert/\|\hat{x}^{(k)}\rVert$ rather than $\kappa(\hat{B}_{1:k})$. We then showed how restarting can be used to stabilize a sketched GMRES implementation, and proposed an adaptive restarting strategy for sGMRES based on our analysis. Our numerical experiments indicate that the convergence of sketched GMRES with our adaptive restarting approach often performs comparably to standard GMRES and has the potential to result in substantial time savings.
@@ Acknowledgments
All authors are supported by the Charles University Research Centre program No. UNCE/24/SCI/005 and by the European Union (ERC, inEXASCALE, 101075632). Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. L. Burke acknowledges funding from the InterFLOP (ANR-20-CE46-0009), NumPEx ExaMA (ANR-22-EXNU-0002), MixHPC (ANR-23-CE46-0005-01), and FPT-4 (ANR-24-CE46- 7572) projects of the French National Agency for Research (ANR).